# Astromer 2

### v1

@@ Abstract
Classical techniques rely on manually engineered features (Debosscher et al., 2007; Nun et al., 2015), which may introduce biases or fail to capture intricate patterns (Pantoja et al., 2022). Foundational models, by processing large volumes of data, have the potential to reveal novel, precise structures in the data. However, this gain in representational power comes at the expense of reduced interpretability.
@@
In this work, Sect. 2 describes the main characteristics of the original version of Astromer, namely Astromer 1. Section 3 introduces the improvements to the model. Sections 4 and 5 describe the data and the main results of this work, respectively. Finally, Sect. 6 outlines the main findings and conclusions.
@@ 3.2 Encoder
The encoder of Astromer 2 has a significantly larger number of parameters, increasing from to . An eight fold increase. This growth is due to the inclusion of six attention blocks, with each block containing four heads and 64 units. Additionally, we have incorporated a dropout layer after the self-attention calculation, as depicted in Fig. 3.
@@ 5 Results
Figure 6 shows the learning curves from the pretraining of Astromer 2. The training took approximately 3 days using 4 A5000 GPUs. The model achieved R2 with a root mean squared error (RMSE) of 0.113 on the probed subset. Astromer 1 had an RMSE of 0.148, making Astromer 2 0.035 better in terms of reconstruction error.
@@ 5.3 Visualizing reconstruction
To gain insight into Astromer’s representations, we visualize the attention values after finetuning. Figure 8 presents two examples showing the mean attention weights from each attention head, along with the mean across all attention heads. For visualization purposes, light curves associated with the average between heads were folded; however, Astromer does not receive folded inputs during processing. We display only the first attention block, as it is more intuitive. This contrasts with intermediate layers, where attention is computed over abstract embeddings.
@@ 5.4 Classification
Figure 11 presents the F1 scores for the Alcock dataset across different scenarios. For comparison, it also includes F1 scores from Astromer 1, evaluated on the same dataset, using both weighted per-sample embeddings (v0) and non-weighted embeddings as detailed in Donoso-Oliva, C. et al. (2023).
The improvements are most pronounced when evaluating classification performance on the ATLAS dataset. As depicted in Fig. 12, the new version of Astromer exhibits a significant advantage in generalizing to other datasets. With just 20 SPC, Astromer achieves a F1-score improvement of over . Thus, Astromer’s performance with 20 SPC surpasses the results previously reported 500 SPC.
@@
In Figures 14 and 15, we present t-SNE projections of the embeddings associated with the first fold of the test set. These projections are computed from the average output of each attention block. Each output consists of a set of vectors, which we reduce to a single vector by averaging.

### v2

# Generalizing Across Astronomical Surveys: Few-Shot Light Curve Classification with Astromer 2

@@ Abstract
We note that self-supervised learning by itself does not necessarily define a foundation model. Foundation models typically involve large-scale pretraining on massive datasets combined with the ability to generalize and adapt effectively across a wide range of downstream tasks, often through fine-tuning. Our approach aligns with this broader concept, but self-supervision is one component within that larger framework
Self-supervised learning enables the use of vast amounts of unlabeled data, significantly increasing the effective training volume beyond what is possible with labeled samples alone. This expanded training capacity allows models to learn richer, more generalizable representations that improve downstream performance even when labeled data is scarce.
Classical techniques rely on manually engineered features (Debosscher et al., 2007; Nun et al., 2015), which may introduce biases or fail to capture intricate patterns (Pantoja et al., 2022). Foundational models, by processing large volumes of data, have the potential to reveal novel, precise structures in the data. However, this gain in representational power comes at the expense of reduced interpretability. While self-supervised deep learning models can offer powerful representations, their interpretability remains limited compared to mechanistic, physics-based models that have clear, physically grounded parameters. However, their interpretability is comparable to other deep learning approaches.
To contextualize our work within the broader field, we reference recent advances in time-series self-supervised learning, such as the Foundation Models for Time Series (Liang et al., 2024), which explore similar concepts in the machine learning community.
@@
The motivation for using self-supervised learning in Astromer is consistent with that of foundational models across AI, including large language models and some of the vision transformers. In the context of astronomical time-series, labeled data are often scarce and costly to obtain at scale. Self-supervised masked modelling enables leveraging vast amounts of unlabeled data to learn powerful, general representations that can be fine-tuned with limited labels for downstream tasks such as classification. This approach has become a central paradigm in modern AI and guides the design of Astromer too.
Our self-supervised masked-modelling approach is inspired by successful techniques in natural language processing (e.g., BERT) and computer vision (e.g., ViTs), which have demonstrated the power of learning rich representations by reconstructing missing parts of the input.
@@
In this work, Section 2 revisits the main characteristics of the original version of Astromer, namely Astromer 1. While a detailed explanation is available in our previous paper, we dedicate a significant portion of this section to reiterating its architecture for several key reasons. Firstly, much of the formulation of Astromer 2 is fundamentally built upon the principles of Astromer 1. Secondly, in response to community questions and to reflect our own improved understanding of its internal processes, we have rewritten the description of the architecture in a clearer and more concise manner. With this enhanced explanation, we aim for this paper to be self-contained, allowing readers to fully comprehend our work without referencing the previous publication. Section 3 then introduces the improvements that define the new model. Sections 4 and 5 describe the data and the main results of this work, respectively. Finally, Section 6 outlines the main findings and conclusions.
@@ 3.2 Encoder
The encoder of Astromer 2 has a significantly larger number of parameters, increasing from to . A six fold increase. This growth is due to the inclusion of six attention blocks, with each block containing four heads and 64 units. Additionally, we have incorporated a dropout layer after the self-attention calculation, as depicted in Fig. 3.
@@ 4.2 Labeled data
The labeled classes in the MACHO dataset were produced through manual labeling supported by statistical analysis and feature extraction techniques, providing a curated and expert-classified dataset. For the ATLAS dataset, the labels are derived from automated classification pipelines based on template fitting and light curve features, which may introduce some uncertainty and biases inherent to automated labeling methods.
@@ 5 Results
Figure 6 shows the learning curves from the pretraining of Astromer 2. The training took approximately 3 days using 4 A5000 GPUs. The model achieved an of 222 (coefficient of determination) measures the proportion of variance in the observed data explained by the model. It ranges from () to 1, with 1 indicating perfect prediction.with a root mean squared error (RMSE) of 0.113 on the probed subset. Astromer 1 had an RMSE of 0.148, making Astromer 2 0.035 better in terms of reconstruction error.
@@ 5.3 Visualizing reconstruction
Figures 8 and 9 show the mean attention weights from each attention head, along with the mean across all heads, for sample light curves from the Alcock and ATLAS datasets, respectively. For each dataset, we present one representative example for each class to illustrate the model’s focus. For visualization purposes, the light curves were folded by their period; however, Astromer does not receive folded inputs during processing. We display the attention from the first attention block, as its interpretation is more intuitive compared to intermediate layers, where attention is computed over more abstract embeddings. As shown in both figures, each attention head learns to focus on different parts of the sequence. Notably, there is a consistent pattern where attention is most strongly concentrated at points of maximum and minimum brightness, suggesting these are identified as key features for reconstruction and characterization. This demonstrates a consistent feature-learning strategy across different datasets and variable star types.
@@ 5.4 Classification
Figure 11 presents the F1 scores for the Alcock dataset across different scenarios. We focus our analysis on the improvements over our previous model, rather than re-evaluating against classical models, as this benchmark was already established in our initial work (Donoso-Oliva, C. et al., 2023). For a direct comparison, the figure includes the F1 scores from Astromer 1, using both weighted per-sample embeddings (A1) and non-weighted embeddings (A2) as detailed in our previous paper.
The improvements are most pronounced when evaluating classification performance on the ATLAS dataset. As depicted in Figure 12, the new version of Astromer exhibits a significant advantage in generalizing to other datasets. With just 20 SPC, Astromer achieves a F1-score improvement of over . Thus, Astromer’s performance with 20 SPC surpasses the results previously reported 500 SPC.
@@
Figures 14 and 15 present t-SNE projections of the test set embeddings, derived from the output of each attention block for the first data fold. Unlike standard text-based BERT models, our architecture does not employ a [CLS] token to obtain a summary representation of the sequence. Therefore, an aggregation step is necessary to produce a single vector for each time series. To achieve this, we reduce the (200, 256) output tensor from each block by averaging across the 200 time steps, resulting in a representative (1, 256) vector.
This averaging strategy was selected over other alternatives, such as concatenation and the maximum, as it consistently produces more coherent clusters for visualization. A detailed quantitative justification supporting this decision is provided in the Appendix A
@@
A key advantage of our self-supervised approach is its ability to enable object retrieval and similarity searches directly in the embedding space without requiring labeled data. As shown in Figures 14 and 15, the model organizes light curves with similar characteristics close to one another, which facilitates effective clustering and retrieval. This capability is particularly valuable in astronomical datasets where labeling is costly and incomplete, providing a powerful tool for exploratory data analysis and discovery beyond classification tasks.
@@ 7 Ethical and Practical Considerations
Appendix A Embedding Visualization Method
To select an optimal method for visualizing the class-based clustering of our pre-trained encoder embeddings, we conducted a quantitative comparison of three distinct aggregation strategies. The methods evaluated were: averaging the token embeddings across the time dimension, concatenating all token embeddings, and applying max pooling across the time dimension.
We used the Silhouette Score with a cosine metric as the evaluation criterion to determine which method produced the most coherent and well-separated clusters. The score for each method was calculated at the output of each of the six attention blocks in the encoder.
The results of this analysis are presented in Figure 18. The key findings from this comparison are:
The averaging method consistently achieves the highest Silhouette Score in the later, more specialized layers of the encoder (blocks 3, 4, and 5) , peaking at a score above . This indicates that averaging produces the most distinct visual clusters.
The max pooling method shows competitive performance in the middle layers but degrades sharply in the final blocks, suggesting it does not capture the final, refined embedding structure as effectively.
The concatenation method results in a score of approximately zero across all blocks, implying it does not produce a useful clustering structure for visualization.
Based on this evidence, we selected the averaging strategy for all embedding visualizations presented in this work.
Appendix B Examples of Masked Value Reconstruction
To provide a qualitative assessment of the model’s ability to learn the underlying structure of the light curves, we present examples of its reconstruction performance on masked input data.
The following figures show a detailed analysis for a selection of variable star classes: Cepheid (Cep_0 , Cep_1 ), RR Lyrae (RRab , RRc ), Eclipsing Binary (EC ), and Long Period Variable (LPV ). Each figure displays a grid corresponding to three randomly selected light curves from its respective class.
Each row in the grid provides three different views of the model’s reconstruction for a single sample:
1.
Left Column (Time Series Reconstruction): This panel displays the original light curve in the time domain. It shows the unmasked observations (Original Visible ), the true values of the masked points (Original Masked ), and the model’s corresponding predictions (Model Prediction ).
2.
Center Column (Prediction vs. True): This panel offers a direct one-to-one comparison of the predicted versus true magnitudes for the masked data points. A dashed line represents a perfect prediction, and the Root Mean Square Error (RMSE) is reported for each sample.
3.
Right Column (Folded Reconstruction): This panel presents the same reconstruction of masked points but in phase space, with the light curve folded by its dominant period. This view is crucial for assessing how well the model captures the periodic nature of the stellar variability.